# Nature Language Model: Deciphering the Language of Nature for Scientific Discovery

### v2

@@ 2.1 Pre-training data
    Linking DNA with Proteins Through the Central Dogma: For DNA sequences, we identify segments that can be transcribed and translated into proteins, following the central dogma of molecular biology. These identified DNA segments are then replaced with the equivalent protein sequences, establishing a direct connection between the genetic blueprint and its functional protein products. This method not only reflects the biological process but also creates a dataset that encapsulates the relationship between nucleotide and amino acid sequences.
@@
The statistics of cross-domain data is in Table 1. Both interleaved sequences and text-science parallel data are types of cross-domain data that aim to facilitate cross-domain interaction. For interleaved sequences, the sources are literature, which can cover a broader range of general topics and wider domains. In contrast, parallel data sources are existing databases that focus on specific properties. Although the topics covered by parallel data are not as diverse as those in interleaved sequences, the amount of data available for each given property is greater. These distinctions highlight the complementary nature of the two types of cross-domain data.
@@ 2.2 Post-training data
We curated a dataset for post-training with about 5.1 million instruction-response pairs encompassing six domains, small molecules, proteins, materials, DNA, RNA and general text (Figure 5). The dataset includes over 60 sub-tasks. For each sub-task, multiple prompts were manually crafted to form diverse instruction-response pairs, covering essential scientific tasks such as molecular optimization, antibody design, and guide RNA design. We provide two examples below:
@@ 2.3 Model architecture
NatureLM models are built upon well-trained large language models (LLMs) with some additional parameters for newly introduce scientific tokens. We used Llama 3 8B dubey2024llama3herdmodels and Mixtral 8x7B jiang2024mixtralexperts to initialize the main part of NatureLM and continued pre-training using the science data described in Section 2.1. Additionally, we trained a model with 1B parameters, which replicates the structural design of Llama 3 but with a reduced number of layers and smaller hidden dimensions. Its pre-training begins with a random selection of 300 billion pure text tokens from the SlimPajama dataset cerebras2023slimpajama , followed by the science data we collected in Section 2.1. This approach ensures a consistent training methodology across all three models. The details of the model architecture are provided in Table 2.
@@ 2.4 Continued pre-training
The validation loss for the three versions of the models is illustrated in Fig. 6. All validation losses decrease as the model size increases from 1 billion to 8 billion, and 8x7 billion. This indicates that larger models are better at capturing the underlying patterns in the data, which is expected due to their increased capacity. The most significant decreases are observed in the text and protein data, suggesting that these datasets benefit more from larger models.
Figure 6: Validation loss for the 1B, 8B, and 8x7B models of NatureLM. Larger models result in smaller validation losses across each domain. NatureLM (8B) is short for “Llama 3 8B NatureLM” throughout this paper.
@@ 2.6 Inference acceleration
As NatureLM will be tested on many downstream tasks, we need to accelerate inference speed to save computational cost. We adopted the following approaches: (1) PagedAttention pagedattention , which optimizes LLM serving by partitioning the key-value (KV) cache into fixed-size, non-contiguous blocks, reducing memory fragmentation and enabling efficient memory sharing; and (2) Selective Batching orca , which batches compatible operations while handling attention separately, allowing for flexible and efficient processing of requests with varying input lengths. We employed the vLLM framework vllm2024 to serve NatureLM models, leveraging its implementations of both PagedAttention and Selective Batching. These optimizations were applied to the 1B, 8B, and 8×7B models. Consequently, the inference speed for the NatureLM 8×7B model reached approximately 525 tokens per second with Brain Float 16 (BF16) precision on two NVIDIA A100 GPUs.
@@ 3.1 Unconditional molecular generation
Table 3: Unconditional evaluation of small molecules generation.
@@ 3.2 Property-to-molecule generation
Figure 7: Evaluation of property-to-molecule generation. (a) Bar plot of the Spearman correlation coefficients between the input property values and generated molecules’ property values. (b) Violin plot showing the input TPSA values and generated molecules’ TPSA values. More properties are left in Fig. LABEL:fig:basic_to_cmpd_violinplot. (c) The joint distribtion of the generated molecules’ TPSA and QED given the input “TPSA=40, QED=0.5” (see Fig. LABEL:fig:qed_fsp3_joint_optim for more cases).
@@ 3.3 Translation between SMILES and IUPAC
We evaluate NatureLM on the translation between SMILES and IUPAC on NC-I2S and NC-S2I yu2024llasmol , the bidirectional IUPAC-SMILES translation dataset comprising 2993 pairs of SMILES and their corresponding IUPAC names (Table 4). We ensure that there is no test set leakage in this setting. On both text-to-SMILES and SMILES-to-text translation tasks, NatureLM (8x7B) outperforms all competing language models in terms of accuracy, demonstrating our model’s strong learning capability for text-molecule correspondence. NatureLM significantly outperforms GPT-4 and Claude 3 Opus claude3 , strong generalist large language models (LLMs), highlighting the necessity of training on scientific data. Compared with another LLM trained on text and SMILES corpus LlaSMolMistral yu2024llasmol , NatureLM also obtains significantly better performance. Moreover, NatureLM (8x7B) performs comparably with STOUT rajan2024stout , the widely-used model trained specially for IUPAC-SMILES translation task, demonstrating NatureLM’s potential as a scientific generalist in specific domains. The performance increases from NatureLM (1B) to NatureLM (8x7B), exhibiting the scaling benefits of larger models. A case study is presented in Fig. S11, comparing NatureLM with general large language models and highlighting the advantages of NatureLM in scientific tasks.
@@
3.4 Target-aware hit generation and optimization
The task is to generate small molecule compounds given the target protein sequence. The combination of NatureLM and structure-based compound design will be explored in the future. We test NatureLM within two distinct scenarios:
@@ 3.4 Target-aware hit generation and optimization
(2) Generate molecular fragments based on the target protein sequences and partial molecular structures as inputs. This method is instrumental during the lead optimization phase, where we scrutinize and refine the molecular architecture to amplify efficacy and precision.
@@
In the first scenario, we compare NatureLM with a sequence generation method, TamGen TamGen , and two approaches that design compounds in 3D space based on the input target: a diffusion-based method, TargetDiff targetdiff , and an autoregressive generation method in 3D space, Pocket2Mol pocket2mol . We follow the evaluation procedure outlined in the TamGen paper TamGen , which includes calculating the docking score using AutoDock Vina, as well as assessing the QED, synthetic accessibility scores (SAS), diversity of the generated compounds, the percentage of compounds with logP in the range [0,5], and the percentage of compounds satisfying the rule-of-five. The results are presented in Table 5. We can see that in terms of docking score, QED and synthesis ability, NatureLM surpasses previous baselines, highlighting its effectiveness.
@@
Additionally, we utilize NatureLM for fragment generation. We selected three papers published after May 2024 Tangallapally2024pdb6PE6 ; Tarr2024pdb9BCG ; Mammoliti2024pdb3xln , where part of their task is to solve the issue of compound optimization. In this context, the input includes a target protein and a backbone that needs optimization. The results are illustrated in Fig. 8. In this instance, it is evident that larger models typically yield superior docking scores.
Figure 8: Violin plot of docking scores for molecules in target-to-fragment generation. This violin plot presents the docking scores of molecules involved in target-to-fragment generation. We selected three recent papers that focus on fragment optimization: Tarr2024pdb9BCG , Mammoliti2024pdb3xln and Tangallapally2024pdb6PE6 , which utilize PDB IDs 9BCG, 3LXN, and 6PE6, respectively. The input fragment is visualized alongside its corresponding PDB ID for clarity.
@@ 3.5 Text-guided binding affinity optimization
To further improve the binding affinity between a target and a molecule, we propose a text-guided binding affinity optimization task. Given a target name and a molecule with a known binding affinity for that target, we aim to generate molecules with higher binding affinity, which is crucial for lead optimization. An example is shown below:
@@
Here, the target information is provided in text format, which complements the FASTA representation used in Section 3.4. We will combine them in the future.
We test NatureLM on 12 targets that are not present in the post-training data and use a hybrid retrieval and docking approach for evaluation. Specifically, for the generated molecules, if we can retrieve their binding affinity values from the ChEMBL database, we compare these values with the original molecule’s binding affinity. Otherwise, we compare their docking scores with the original molecule. For the 12 selected targets, their Spearman correlation between the docking score and the actual binding affinity for known molecules exceeds 0.5, indicating the reliability of using docking for assessment (Table S4).
We observe that NatureLM can successfully improve the molecule’s binding affinity by making small modifications on its chemical components, much like what a chemist would typically do (Fig. 9b). Compared with GPT-4, NatureLM can generate more molecules (Fig. 9a) with higher binding affinity, making it a better tool for molecule optimization than general domain LLM. Another observation is that more than 90% molecules generated by NatureLM do not have known binding affinity score in ChEMBL database. For 8 out of the 12 targets, over 50% of the generated novel molecules successfully decreased the docking scores (Fig. S5), demonstrating the model’s potential in exploring chemical spaces and discovering new drugs. We observe that NatureLM (8x7B) and NatureLM (8B) outperform NatureLM (1B) as they generate more correct molecules for the majority of targets (Fig. S6).
@@ 3.6 Text-guided metabolism & distributional property optimization
Table 6 displays the outcomes of BBBP and metabolism optimization. The success rates for optimizing BBBP with the 1B, 8B, and 8x7B versions of NatureLM are 0.482, 0.549, and 0.552, respectively. Larger models show better performance, though the improvement is not substantial. This suggests potential for enhancement opportunities in the future. For metabolism optimization, generally, the 8B model outperforms the others in terms of success rate, followed by the 8x7B model and lastly the 1B model. The 1B and 8B models share the same architecture (dense models, large vocabulary size), whereas the 8x7B model has a distinct one (mixture-of-expert model, relative small vocabulary size). In this particular task, the progression from the 1B model to the 8B model is consistent. However, a detailed analysis contrasting the 8x7B model is to be conducted in subsequent studies. Additionally, we jointly optimized metabolism and a basic property. The findings indicate that larger models generally yield better results (see Table S5).
@@ 3.7 Retrosynthesis prediction
We followed the common practice for splitting the USPTO-50K dataset dai2019retrosynthesis ; maziarz2024re , and evaluated the performance using the 5007 reactions included in the test set. We ensured that there is no test set leakage in this setting. As outlined in Table 7, all sizes of NatureLM models surpass other methods in terms of top-$k$ accuracy, demonstrating our model’s accurate predictive ability for retrosynthesis prediction. NatureLM significantly outperforms GPT-4, a general LLM trained on human languages. This suggests that training on scientific data is crucial for models to excel in scientific tasks. Furthermore, NatureLM outperforms the state-of-the-art domain-specific models such as LocalRetro chen2021localretro and R-SMILES Zhong2022rsmiles , showing NatureLM’s potential as a scientific generalist in critical scientific tasks. We also note an increase in performance from NatureLM (1B) to NatureLM (8x7B), demonstrating the scaling advantages of larger models.
@@
We selected a reaction from a U.S. patent with ID US12018024B2 (granted to Novartis on June 25, 2024) as the case study. As shown in Figure 10, the product is a brominated heterocyclic compound. To synthesize such molecules, organic chemists typically begin by constructing the ring system, then followed by a halogenation reaction on the ring. The halogenated site is subsequently used in further syntheses, often coupling with other ring systems. In this context, our NatureLM model accurately predicted one of the most common brominating agent NBS (N-Bromosuccinimide) in this step, aligning perfectly with experimental results. DFT optimization calculations to reactant CCOC(=O)c1cc2ccnn2cn1 show that 3-position is the most nucleophilic. So bromination prefers at 3-position. This demonstrates NatureLM’s capability to effectively predict useful reactants in chemical reactions.
In contrast, DeepSeek-R1 deepseekai2025r1 model selected incorrect reactants. DeepSeek-R1 focused on constructing a heterocyclic ring system in this step. However, it incorrectly positioned the nitrogen and carbon atoms in the substrate, leading to the wrong outcome. Meanwhile, o3-mini-high model selected another correct route but not so convenient for the whole synthetic process. o3-mini-high might attempt to construct the acetyl side chain by transition-metal-catalyzed CO insertion reaction. Although radical mechanism can achieve this since iodo atom is more reactive than bromo atom, introducing the side chain at this stage may still lead to a side reaction at the bromo atom. In the actual synthesis, this acetyl side chain is generally constructed in previous steps. These cases illustrate that, so far, general LLMs cannot fully grasp the rationale behind chemical synthesis. Although they can generate some reasons and predictions, these do not align with common chemist strategies.

### v3

@@ 2.1 Pre-training data
    Linking DNA with Proteins Through the Central Dogma: For DNA sequences, we identify segments that can be transcribed and translated into proteins, following the central dogma of molecular biology. These identified DNA segments are then replaced with the equivalent protein sequences, establishing a direct connection between the genetic blueprint and its functional protein products. This method not only reflects the biological process but also creates a dataset that encapsulates the relationship between nucleotide and amino acid sequences. We retrieved data from the RefSeq database [^9] https://www.ncbi.nlm.nih.gov/refseq/ and extracted protein sequences, including their isoforms, from annotated genes, along with their corresponding flanking DNA sequences.
@@
The statistics of cross-domain data is in Table 1. Both interleaved sequences and text-science parallel data are types of cross-domain data that aim to facilitate cross-domain connections. For interleaved sequences, the sources are literature, which covers a broader range of general topics and wider domains. In contrast, parallel data sources are existing databases that focus on specific properties. Although the topics covered by parallel data are not as diverse as those in interleaved sequences, the amount of data available for each given property is larger. These distinctions highlight the complementary nature of the two types of cross-domain data.
@@ 2.2 Post-training data
We curated a dataset for post-training with about 5.1 million instruction-response pairs encompassing six domains, small molecules, proteins, materials, DNA, RNA and general text (Figure 5). The dataset includes over 60 categories of tasks. For each task category, multiple prompts were manually crafted to form diverse instruction-response pairs, covering essential scientific tasks such as molecular optimization, antibody design, and guide RNA design. We provide two examples below:
@@ 2.3 Model architecture
NatureLM models are built upon well-trained large language models (LLMs) with some additional parameters for newly introduce scientific tokens. We used Llama 3 8B dubey2024llama3herdmodels and Mixtral 8x7B jiang2024mixtralexperts to initialize the main part of NatureLM and continued pre-training using the science data described in Section 2.1. Additionally, we trained a model with 1B parameters, which replicates the structural design of Llama 3 but with a reduced number of layers and smaller hidden dimensions. The pre-training of NatureLM 1B begins with a random selection of 300 billion pure text tokens from the SlimPajama dataset cerebras2023slimpajama , followed by the science data we collected in Section 2.1. This approach ensures a consistent training methodology across all three models. The details of the model architecture are provided in Table 2.
@@ 2.4 Continued pre-training
The validation loss for the three versions of the models is illustrated in Fig. 6. All validation losses decrease as the model size increases. This indicates that larger models are better at capturing the underlying patterns or rules in the data, which is expected due to their increased capacity. The most significant decreases are observed in the text and protein data, suggesting that these datasets benefit more from enlarged models.
Figure 6: Validation loss for the 1B, 8B, and 8x7B NatureLM models. Larger models result in smaller validation losses across all domains. NatureLM (8B) is short for “Llama 3 8B NatureLM” throughout this paper.
@@ 2.6 Inference acceleration
As NatureLM will be tested on many downstream tasks, we need to accelerate inference speed to reduce computational cost. We adopted the following approaches: (1) PagedAttention pagedattention , which optimizes LLM serving by partitioning the key-value (KV) cache into fixed-size, non-contiguous blocks, reducing memory fragmentation and enabling efficient memory sharing; and (2) Selective Batching orca , which batches compatible operations while handling attention separately, allowing for flexible and efficient processing of requests with varying input lengths. We employed the vLLM framework vllm2024 to serve NatureLM models, leveraging its implementations of both PagedAttention and Selective Batching. These optimizations were applied to the 1B, 8B, and 8×7B models. Consequently, the inference speed for the NatureLM 8×7B model reached approximately 525 tokens per second with Brain Float 16 (BF16) precision on two NVIDIA A100 GPUs.
@@ 3.1 Unconditional molecular generation
Table 3: Unconditional evaluation of small molecules generation. RDKit is used to convert the generated SMILES strings into molecular structures and check validity. The uniqueness ratio is calculated among the valid molecules.
@@ 3.2 Property-to-molecule generation
Figure 7: Evaluation of property-to-molecule generation. (a) Bar plot of the Spearman correlation coefficients between the input property values and generated molecules’ property values. (b) Violin plot showing the input TPSA values and generated molecules’ TPSA values. More properties are shown in Fig. LABEL:fig:basic_to_cmpd_violinplot. (c) The joint distribtion of the generated molecules’ TPSA and QED given the input “TPSA=40, QED=0.5” (see Fig. LABEL:fig:qed_fsp3_joint_optim for more cases).
@@ 3.3 Translation between SMILES and IUPAC
We evaluate NatureLM on the translation between SMILES and IUPAC on NC-I2S and NC-S2I yu2024llasmol , the bidirectional IUPAC-SMILES translation dataset comprising 2993 pairs of SMILES and their corresponding IUPAC names (Table 4). We ensure that there is no test set leakage in this setting. On both text-to-SMILES and SMILES-to-text translation tasks, NatureLM (8x7B) outperforms all competing language models in terms of accuracy, demonstrating our model’s strong capability for text-molecule correspondence. NatureLM significantly outperforms GPT-4 and Claude 3 Opus claude3 , strong generalist large language models (LLMs), highlighting the necessity of training on scientific data. Compared with another LLM trained on text and SMILES corpus LlaSMolMistral yu2024llasmol , NatureLM also obtains significantly better performance. Moreover, NatureLM (8x7B) performs comparably with STOUT rajan2024stout , the widely-used model trained specially for IUPAC-SMILES translation task, demonstrating NatureLM’s potential as a scientific generalist in specific domains. The performance increases from NatureLM (1B) to NatureLM (8x7B), exhibiting the scaling benefits of larger models. A case study is presented in Fig. S11, comparing NatureLM with general large language models and highlighting the advantages of NatureLM in scientific tasks.
@@
3.4 Target-aware hit compound generation and optimization
The task is to generate small molecule compounds given the target protein sequence. The combination of NatureLM and structure-based compound design, such as TamGen TamGen and TargetDiff targetdiff , will be explored in the future. We test NatureLM within two distinct scenarios:
@@ 3.4 Target-aware hit compound generation and optimization
(2) Generate molecular fragments based on the target protein sequences and partial molecular structures as inputs. This method is instrumental during the lead optimization phase, where we scrutinize and refine the molecular architecture to improve efficacy and precision.
@@
In the first scenario, we compare NatureLM with a sequence generation method, TamGen TamGen , and two other approaches that design compounds in 3D space based on the input target structure: a diffusion-based method, TargetDiff targetdiff , and an autoregressive generation method in 3D space, Pocket2Mol pocket2mol . We follow the evaluation procedure outlined in the TamGen paper TamGen , which includes calculating the docking score using AutoDock Vina autodock_vina_1 ; autodock_vina_2 , as well as assessing the QED, synthetic accessibility scores (SAS), diversity of the generated compounds, the percentage of compounds with logP in the range [0,5], and the percentage of compounds satisfying the rule-of-five. The results are presented in Table 5. We can see that in terms of docking score, QED and synthesis ability, NatureLM surpasses previous baselines, highlighting its effectiveness. For the other metrics, NatureLM achieves comparable performance to the baselines.
@@
Additionally, we utilize NatureLM for fragment generation. We selected three papers published after May 2024 Tangallapally2024pdb6PE6 ; Tarr2024pdb9BCG ; Mammoliti2024pdb3xln , where part of their task is to solve the issue of compound optimization. From these papers, we choose the target proteins and the fragments to retain, allowing NatureLM to generate the remaining fragments. The results are illustrated in Fig. 8. In this instance, it is evident that larger NatureLM models yield superior docking scores in general.
Figure 8: Docking scores for molecules in target-to-fragment generation. This violin plot presents the docking scores of molecules involved in target-to-fragment generation. We selected three recent papers that focus on fragment optimization: Tarr2024pdb9BCG , Mammoliti2024pdb3xln and Tangallapally2024pdb6PE6 , which utilize PDB IDs 9BCG, 3LXN, and 6PE6, respectively. The input fragment is visualized alongside its corresponding PDB ID for clarity.
@@ 3.5 Text-guided binding affinity optimization
To further improve the binding affinity between a target and a molecule, we propose a text-guided binding affinity optimization task. Given a target name and a molecule with a known binding affinity for that target, we aim to generate molecules with higher binding affinity, which is a crucial component for lead optimization. An example is shown below:
@@
Here, the target information is provided in text format, which complements the FASTA representation used in Section 3.4. We will combine text and FASTA in the future.
We test NatureLM on 12 targets that are not present in the post-training data and use a hybrid retrieval and docking approach for evaluation. Specifically, for the generated molecules, if we can retrieve their binding affinity values from the ChEMBL database, we compare these values with the original molecule’s binding affinity. Otherwise, we resort to docking scores as the comparison metric. For the 12 selected targets, their Spearman correlation between the docking score and the actual binding affinity for known molecules exceeds 0.5, indicating the reliability of using docking for assessment (Table S4).
We observe that NatureLM can successfully improve the molecule’s binding affinity by making modifications on its chemical components, in a manner similar to what a chemist would typically do (Fig. 9b). Compared with GPT-4, NatureLM can generate molecules (Fig. 9a) with higher binding affinity, making it a better tool for molecule optimization than general purpose LLM. Another observation is that more than 90% molecules generated by NatureLM do not have known binding affinity score in ChEMBL database. For 8 out of the 12 targets, over 50% of the generated novel molecules successfully improved the docking scores (Fig. S5), demonstrating the model’s potential in exploring chemical spaces and discovering novel drugs. We observe that NatureLM (8x7B) and NatureLM (8B) outperform NatureLM (1B) as they generate more correct molecules for the majority of targets (Fig. S6).
@@ 3.6 Text-guided metabolism & distributional property optimization
Table 6 displays the outcomes of BBBP and metabolism optimization. The success rates for optimizing BBBP with the 1B, 8B, and 8x7B versions of NatureLM are 0.482, 0.549, and 0.552, respectively. Larger models show better performance, suggesting potential for enhancement opportunities in the future. For metabolism optimization, generally, the 8B model outperforms the others in terms of success rate, followed by the 8x7B model and lastly the 1B model. The 1B and 8B models share the same architecture (dense models, large vocabulary size), whereas the 8x7B model has a distinct one (mixture-of-expert model, relative small vocabulary size). In this particular task, the progression from the 1B model to the 8B model is consistent. However, a detailed analysis contrasting the 8x7B model is to be conducted in subsequent studies. Additionally, we jointly optimized metabolism and a basic property. The findings indicate that larger models generally yield better results (see Table S5).
@@ 3.7 Retrosynthesis prediction
We followed the common practice for splitting the USPTO-50K dataset dai2019retrosynthesis ; maziarz2024re , and evaluated the performance using the 5007 reactions included in the test set. We ensured that there is no test set leakage in this setting. As outlined in Table 7, all sizes of NatureLM models surpass other methods in terms of top-$k$ accuracy, demonstrating our model’s accurate predictive ability for retrosynthesis prediction. NatureLM significantly outperforms GPT-4, suggesting that training on scientific data is crucial for models to excel in scientific tasks. Furthermore, NatureLM outperforms the state-of-the-art domain-specific models such as LocalRetro chen2021localretro and R-SMILES Zhong2022rsmiles , showing NatureLM’s potential as a scientific generalist in critical scientific tasks. We also note an increase in performance from NatureLM (1B) to NatureLM (8x7B), demonstrating the scaling advantages of larger models.
@@
We selected a newly discovered reaction from a U.S. patent with ID US12018024B2 (granted to Novartis on June 25, 2024) as the case study. As shown in Figure 10, the product is a brominated heterocyclic compound. To synthesize such molecules, organic chemists typically begin by constructing the ring system, then followed by a halogenation reaction on the ring. The halogenated site is subsequently used in further syntheses, often coupling with other ring systems. In this context, our NatureLM model accurately predicted one of the most common brominating agent NBS (N-Bromosuccinimide) in this step, aligning perfectly with experimental results. DFT optimization calculations to reactant CCOC(=O)c1cc2ccnn2cn1 show that 3-position is the most nucleophilic. So bromination prefers at 3-position. This demonstrates NatureLM’s capability to effectively predict useful reactants in chemical reactions.
In contrast, DeepSeek-R1 deepseekai2025r1 model selected incorrect reactants. DeepSeek-R1 focused on constructing a heterocyclic ring system in this step. However, it incorrectly positioned the nitrogen and carbon atoms in the substrate, leading to the wrong outcome. Meanwhile, o3-mini-high model selected another correct route but not so convenient for the whole synthetic process. o3-mini-high might attempt to construct the acetyl side chain by transition-metal-catalyzed CO insertion reaction. Although radical mechanism can achieve this since iodo atom is more reactive than bromo atom, introducing the side chain at this stage may still lead to a side reaction at the bromo atom. In the actual synthesis, this acetyl side chain is generally constructed in previous steps. These cases illustrate that, so far, general LLMs cannot fully grasp the rationale behind chemical synthesis. Although they can generate some reasons and predictions, these do not align with practical strategies.